A box contains 5 red balls, 4 blue balls, and 3 green balls. Two balls are drawn one after another without replacement. What is the probability that both balls are of different colors?
- AOption (a)
18/29
- BOption (b)
22/29
- COption (c)
17/29
- DOption (d)
47/66
- EOption (e)
7/13
View answer & explanationHide answer & explanation
Correct answer
47/66
Why this answer is correct
Shortcut Trick
Total balls = 5 + 4 + 3 = 12
P(Different colors) = 1 − P(Same color)
P(Same color) = [5C2 + 4C2 + 3C2] ÷ 12C2
P(Same color) = [10 + 6 + 3] ÷ 66 = 19/66
P(Different colors) = 1 − 19/66 = 47/66
∴ The correct answer is 47/66.
Alternate Method
Given: Red balls = 5, Blue balls = 4, Green balls = 3. Total balls = 12. Two balls drawn without replacement.
Formula Used: Probability = (Favorable Outcomes) ÷ (Total Outcomes)
⇒ Total ways to draw 2 balls from 12 = 12C2 = (12 × 11) ÷ (2 × 1) = 66
⇒ Ways to draw 2 balls of same color = (2 Red) OR (2 Blue) OR (2 Green)
⇒ Ways = 5C2 + 4C2 + 3C2
⇒ Ways = [(5×4)/2] + [(4×3)/2] + [(3×2)/2] = 10 + 6 + 3 = 19
⇒ Probability (Same color) = 19/66
⇒ Probability (Different colors) = 1 − 19/66 = 47/66
∴ The correct answer is 47/66.
Additional Information
Complementary Events
The sum of the probability of an event happening and not happening is always 1. P(A) + P(A') = 1.
Combinations Formula
Used for selection where order doesn't matter: nCr = n! ÷ [r! × (n−r)!].
Without Replacement Rule
When drawing items without replacement, the total number of available outcomes decreases by 1 after each draw.